Question # 2. (2 marks) Show that the ratio test fails to apply to the series, 7-n+(-1), but that the root test does apply.
Answer #1

7-n+lim nel = 7-ntion anti= 7-n-1-(4) wing Ratio Test: lim = lim anti int-cuin zon + Gin an nto م = 7-3 even = lim 7-1-26-17

Using root-Test : ac L=im nto = lim 17-ntuinlyn - Mindez+++yn 17-1+01= root Test Series converges. I <I ::. Byn-D f(x) = Ź an (x+1)? +(-4)= El an (-3) diverger. neo 0 +(2) Σ αη 31 → convergen. Radius of converges = 3. lim anal-3n+1 ny